Let CP represent the event that the ankle was truly fractured, and let T+ represent the...

  1. Let CP represent the event that the ankle was truly fractured, and let T+ represent the event that the test said it was. From the table, estimate P(CP|T+) (the positive predictive value and P(CA|T-) (the negative predictive value).

Predicted by the Ottawa Ankle Test Truth Regarding Fracture

Not Fractured | Fractured | Totals

Not Fractured | 51 | 5 | 56

Fractured | 277 | 88 | 365

Totals | 328 | 93 | 421
Answer #1

P(CP|T+) =P(really fractured given tested positive for fracture)=5/93=0.053763

P(CA|T-)=P(not fractured given tested negative for fracture)=277/328=0.844512